OpenStudy (waheguru):

How do I Factor this 2(a+b)^2+5)(a+b)+3

OpenStudy (waheguru):

\[2(a+b)^2+5)(a+b)+3 factor\] Can someone please explain how to factor this?

OpenStudy (mertsj):

Replace (a+b) with x and try that

OpenStudy (waheguru):

\[2(a+b)^2+5(a+b)+3 \]

OpenStudy (mertsj):

\[2x^2+5x+3\]

OpenStudy (anonymous):

Yeah. so let (a+b) = x, Then 2x^2+5x+3

OpenStudy (waheguru):

Why do I replace with ?

OpenStudy (mertsj):

So its easier to factor.

OpenStudy (mertsj):

More familiar looking

OpenStudy (anonymous):

u replace it with X to make the expression easier. Later u can use this x = a+b to find proper value

OpenStudy (waheguru):

Can you also explqin how to solve without replacing it with x

OpenStudy (mertsj):

\[2x^2+5x+3=(2x+3)(x+1)\]

OpenStudy (mertsj):

\[(2x+3)(x+1)=[(2(a+b)+3][(a+b)+1]=\]

OpenStudy (mertsj):

\[(2a+2b+3)(a+b+1)\]

OpenStudy (anonymous):

well uh without replacing with x it becomes very tedious. thats why we do it using x = a +b

OpenStudy (waheguru):

the answer is somthine else that what I am getting? http://www.wolframalpha.com/input/?i=2%28a%2Bb%29%5E2%2B5%28a%2Bb%29%2B3+factor

OpenStudy (anonymous):

no LOL its the same :)

OpenStudy (waheguru):

I do not get how its the same?

OpenStudy (mertsj):

Yep. Thanks to Wolf for proving me right.

OpenStudy (mertsj):

(2a+2b+3)(a+b+1) is the same as (a+b+1)(2a+2b+3)

OpenStudy (mertsj):

4 x 3 = 3 x 4

OpenStudy (anonymous):

he means; 3 x 2 = 6 and even 2 x 3 = 6. This is called as commutative property of multiplication. a x b = b x a.

OpenStudy (waheguru):

(2x+3)(x+1) what did u do after I do not understand

OpenStudy (mertsj):

Replaced x with a+b

OpenStudy (mertsj):

Remember: At the beginning we said we would replace a+b with x so it would be easier to factor

OpenStudy (waheguru):

yea..

OpenStudy (mertsj):

Then we factored it with x and said , oh wait...x is really a+b

OpenStudy (mertsj):

I will leave now. I think part of your confusion is that there is one too many cooks and it is confusing you.

OpenStudy (waheguru):

\[2x2+5x+3=(2x+3)(x+1)\] Should it have been (2x+5)(x+3)

OpenStudy (waheguru):

@Mertsj ?

OpenStudy (agent0smith):

No, the first one was correct.

OpenStudy (mertsj):

Is 5 x 3 = 3???

OpenStudy (waheguru):

becuse we factor out x x(2x+5)+3 and becomes (2x+5)(x+3)

OpenStudy (mertsj):

Ok. You can say those are the factors and I will say that the factors of 3 are 5 x3. When and if you prove me wrong, I will prove you wrong.

OpenStudy (waheguru):

I am not contradicting what you said nor trying to prove you wrong in any way. I just don't understand how you have factored it because I am getting a different answer. I want to learn and that's why questioned you...

OpenStudy (mertsj):

That was not the point. i was hoping you would say "The factors of 3 are not 3x 5 because 3 x 5 = 15 and not 3".

OpenStudy (mertsj):

I was trying to get you to see that whatever you say the factors are must multiply back to the original thing you factored. And the factors of 2x^2+5x+3 could NOT be (2x+5)(x+3) because: